OpenStudy (anonymous):

Please sketch the region of integration and evaluate the integral (see question).

OpenStudy (anonymous):

\[\int\limits_{1}^{4}\int\limits_{0}^{\sqrt{x}}(3/2)e ^{y/\sqrt{x}}dydx\]

sam (.sam.):

\[\huge \int\limits\limits_1^4 \left(\int\limits\limits_0^{\sqrt{x}} \frac{3}{2} e^{\frac{y}{\sqrt{x}}} \, dy\right) \, dx\]

OpenStudy (dumbcow):

well for inside integral, make substitution u = y/sqrtx du = dy/sqrtx \[\rightarrow \frac{3}{2}\sqrt{x}\int\limits_{0}^{\sqrt{x}}e^{u} du = \frac{3}{2}\sqrt{x}(e-1)\] then integrate that wrt x \[ \frac{3}{2}(e-1)\int\limits_{0}^{4}\sqrt{x} dx = (e-1)*4^{3/2} = 8(e-1)\]

OpenStudy (anonymous):

That looks right and is was I got too. Do you know how I should sketch the region of integration?

OpenStudy (dumbcow):

oops i made mistake on limits, i put 0 to 4 instead of 1 to 4

OpenStudy (dumbcow):

which changes answer to 7(e-1)

OpenStudy (anonymous):

Oh right

OpenStudy (dumbcow):

 this would be the region in xy plane